1. In a mass spectrum, aldehydes containing a hydrogen atom in the gamma position will likely exhibit which of the following common fragmentation patterns? Select all that apply. A : McLafferty rearrangement B : dehydration C : beta cleavage D : delta cleavage

2. A molecule containing which of the following atoms will produce M+• and (M+2)+• peaks of equal intensity? A : Nitrogen B : Chlorine C : Oxygen D : Carbon E : Bromine

3. How many rings are found in an alkane with a formula of C 8H 12.
(Enter a whole number only)               
Answer #1

A: Mchafferty rearrangement Ans: E (bromine) OBE = 8-12 a = 9-6=3 T 3 rings